Carol is twice as old as Jane. Five years ago she was 3 times as old as Jane. How old is Jane?
Answer by Maths68(1474)

Let
Present age of Carol = c
Present age of Jane = j

Given
Carol is twice as old as Jane.
c=2j..................(1)
Five years ago
Carol was = c-5
Jane was = j-5
Then
Carol was 3 times as old as Jane.
c-5=3(j-5)
c-5=3j-15
c=3j-15+5
c=3j-10...............(2)
Put the value of c from (1) to (2)
2j=3j-10
2j-3j=-10
-j=-10
Multiply by -1 both sides of above equation
j=10
Put the value of j in (1)
c=2j
c=2(10)
c=20



Present age of Carol = c = 20
Present age of Jane = j = 10